### v1

# Thermodynamic Entropic Uncertainty Relation

Abstract
Thermodynamic uncertainty relations reveal a fundamental trade-off between the precision of trajectory observables and entropy production, where the uncertainty in observables is quantified by their variance. In the context of information theory, uncertainty is often evaluated in terms of Shannon entropy, but it remains unclear whether there is a quantitative relation between Shannon entropy of the observables and entropy production in stochastic thermodynamics. In this Letter, we show that uncertainty relations can be formulated with observable Shannon entropy and entropy production. We introduce symmetry entropy, an entropy measure that quantifies the symmetry of the observable distribution, and demonstrate that a greater asymmetry in the observable distribution demands higher entropy production. This Letter elucidates the role of Shannon entropy of observables within stochastic thermodynamics, thereby establishing a foundation for deriving new uncertainty relations.
Introduction.—Quantum mechanics operates in ways that are fundamentally different from classical physics. The Heisenberg uncertainty relation, proposed by Heisenberg in 1927 [1], captures the unique nature of quantum mechanics through a single inequality, rendering the inability to determine position and momentum precisely. Robertson [2] generalized the Heisenberg uncertainty relation so that the relation can incorporate observables other than position or momentum. Given a quantum state |$\ket{\psi}$ and observables $A$ and $B$, the Robertson uncertainty relation states
$\mathrm{Var}[A]\mathrm{Var}[B]\geq\frac{1}{4}|\braket{\psi}{[A,B]}{\psi}|^{2},$ (1)
where [,][$\bullet$,$\bullet$][ ∙ , ∙ ] is the commutator and $\mathrm{Var}[A]$ is the variance of $A$ with respect to |$\ket{\psi}$. The Robertson uncertainty relation is recognized for its fundamental role in deriving various other relations, including the quantum speed limit [3, 4, 5]. Although the Robertson uncertainty relation given by Eq. (1) quantifies the uncertainty of the observables via their variance, uncertainty is often evaluated using the Shannon entropy in the context of information theory. It is therefore natural to expect uncertainty relations that incorporate the Shannon entropy. Indeed, Ref. [6] showed that the uncertainty relation involving the entropy of position and momentum holds in quantum mechanics. Several generalizations and extensions have been proposed for the entropic uncertainty relation [7] and the most well known instance is Maassen-Uffink relation [8, 9]:
$H[A]+H[B]\geq\ln\frac{1}{c}.$ (2)
Here, $H[A]$ and $H[B]$ denotes the Shannon entropy of measurement outputs of $A$ and $B$, respectively, and $c$ denotes the maximum overlap between two eigenvectors of $A$ and $B$. The entropic uncertainty relation is crucial not only for understanding the nature of quantum mechanics but also plays a vital role in quantum cryptography, especially in quantum key distribution protocols.
In recent years, it has become clear that uncertainty relations are prevalent in stochastic thermodynamics [10, 11]. In particular, the thermodynamic uncertainty relation [12, 13] indicates a trade-off between entropy production and the relative variance of the current observable. For a stochastic thermodynamic system in the steady state and a trajectory observable $F$, the following relation holds:
$\frac{\mathrm{Var}[F]}{\mathbb{E}[F]^{2}}\geq\frac{2}{\Sigma},$ (3)
where $\mathbb{E}[F]$ and $\mathrm{Var}[F]$ denote the expectation and the variance of $F$, respectively, and $\Sigma$ is the entropy production within the time interval of interest. Equation (3) suggests that achieving greater precision requires increased entropy production and signifies a no-free lunch in thermodynamic systems. The thermodynamic uncertainty relation in Eq. (3) resembles the Robertson uncertainty relation in that the uncertainty is evaluated with the variance. In fact, certain types of thermodynamic uncertainty relations are known to actually be derived from the Robertson uncertainty relation [14, 15]. Given this context, a simple question arises: Does an entropic uncertainty relation hold in stochastic thermodynamics? In this paper, we confirm that this conjecture is correct and demonstrate that there is an uncertainty relation between entropy production and the symmetry entropy [cf. Eq. (12)], which quantifies the extent of symmetry of observable distributions. Specifically, we show that the sum of the entropy production and the symmetry entropy should be no less than $\ln 2$ [cf. Eqs. (14) and (16)]. In other words, the asymmetry of the observable probability distribution, as quantified by entropy, requires that the entropy production be at least equal to this measure of asymmetry. We validate the uncertainty relation with a biased random walk, which is often used to model biomolecular processes. The trade-off relationship in thermodynamic cost has been extensively studied, particularly in terms of thermodynamic uncertainty relations [12, 13, 16, 17, 18, 19, 20, 21, 22, 23] using the variance of observable quantities and the speed limits using the distance between states [24, 25, 26, 22, 27, 28, 29]. This study demonstrates a trade-off between the entropy of observables and the entropy production, which is expected to lead to the derivation of other trade-off relations.
Figure 1: Examples of the symmetry entropy $\Lambda[P(\Phi)]$. Horizontal axes denote values of $\Phi$ (left column) and $|\Phi|$ (right column). The values of $\Phi$ always form pairs; $-1$ and $1$, and $-2$ and $2$ are two pairs in the examples. Vertical axes denote probability distribution $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column). (a) Symmetric distribution where probabilities of all of the pairs are identical. $P(|\Phi|)$ is different from the original distribution $P(\Phi)$ and thus $\Lambda[P(\Phi)]$ is $\ln 2$. (b) Asymmetric distribution where probabilities of all of the pairs are fully biased. $P(|\Phi|)$ and $P(\Phi)$ are identical and thus $\Lambda[P(\Phi)]$ is 00. (c) Asymmetric distribution where probabilities of all of the pairs are fully biased. The distributions $P(|\Phi|)$ and $P(\Phi)$ are different, but they effectively become the same if the labels for the pair $-1$ and $1$ are swapped. Therefore, they are essentially the same distribution. This results in $\Lambda[P(\Phi)]=0$.
Methods.—Let $X$ be a random variable and $P(X)$ be its probability distribution. The Shannon entropy $H[P(X)]$ is defined by
$H[P(X)]\equiv-\sum_{x}P(X=x)\ln P(X=x).$ (4)
Let $Y$ be another random variable. The Kullback-Leibler divergence between $P(X)$ and $P(Y)$ is defined by
$D[P(X)\|P(Y)]$ $\equiv\sum_{x}P(X=x)\ln\frac{P(X=x)}{P(Y=x)}$
$=-H[P(X)]+C[P(X),P(Y)].$ (5)
where $C[P(X),P(Y)]$ is the cross entropy:
$C[P(X),P(Y)]\equiv-\sum_{x}P(X=x)\ln P(Y=x).$ (6)
It is known that the Kullback-Leibler divergence is non-negative. Moreover, the Kullback-Leibler divergence satisfies monotonicity. Consider a transformation that maps the original random variables $X$ and $Y$ to new random variables $\tilde{X}$ and $\tilde{Y}$, respectively. Then the following monotonicity relation holds:
$D[P(X)\|P(Y)]\geq D[P(\tilde{X})\|P(\tilde{Y})],$ (7)
where $P(\tilde{X})$ and $P(\tilde{Y})$ are probability distributions of the transformed variables $\tilde{X}$ and $\tilde{Y}$. The Kullback-Leibler divergence is a measure that quantifies the distance between two probability distributions. The monotonicity shown by Eq. (7) implies that no matter what transformation is applied to random variables, the distance between the two probability distributions cannot be increased.
Having introduced basic concepts of the divergence, we move to consideration of stochastic thermodynamic systems. Stochastic thermodynamics considers processes whose state change is described by a stochastic process. Let $\Gamma$ be a stochastic trajectory of the process and $\Gamma^{\dagger}$ be its time reversal. Moreover, we can define the probability of measuring $\Gamma$, which is denoted by $\mathcal{P}(\Gamma)$. Assuming the local detailed balance, it is known that the entropy production under the steady-state condition is defined by the Kullback-Leibler divergence:
$\Sigma=D[\mathcal{P}(\Gamma)\|\mathcal{P}(\Gamma^{\dagger})].$ (8)
Note that the expression of Eq. (8) also holds for continuous processes such as Langevin dynamics. Consider an observable $\Phi(\Gamma)$, which is a function of the trajectory $\Gamma$. Here, we assume that $\Phi(\Gamma)$ is anti-symmetric under the time reversal:
$\Phi(\Gamma)=-\Phi(\Gamma^{\dagger}).$ (9)
For example, $\Phi(\Gamma)$ represents a thermodynamic current. Important thermodynamic quantities, such as stochastic dissipated heat or displacement, are expressed by $\Phi(\Gamma)$. Here, we initially assume that $\Phi(\Gamma)$ takes values in a countable set, that is, the probability distribution $P(\Phi)$ is discrete. However, all the results below hold for the continuous case as well.
Figure 2: Illustration of the example. (a) Biased random walk with right and left transition probabilities are $w_{+}$ and $w_{-}$, respectively. (b) Illustration of F1-ATPase. This biased random walk can model molecular motors such as F1-ATPase, where the motor uses ATP hydrolysis to rotate discretely in steps of $32\pi/$ 3 in the counter-clockwise direction.
Results.—We derive the thermodynamic entropic uncertainty relation, which is the main result of this Letter. Using the monotonicity of Kullback-Leibler divergence [Eq. (7)] and the time reversal property of $\Phi$ [Eq. (9)], we have
$\Sigma$ $\geq D[P(\Phi)\|P(-\Phi)]$
$=-H[P(\Phi)]+C[P(\Phi),P(-\Phi)].$ (10)
Here, the cross entropy is evaluated as $C[P(\Phi),P(-\Phi)]=-\sum_{\phi}P(\Phi=\phi)\ln P(\Phi=-\phi)$. Since the cross entropy term is no less than the entropy, it is non-negative, $C[P(\Phi),P(-\Phi)]\geq H[P(\Phi)]\geq 0$. An important observable is that $C[P(\Phi),P(-\Phi)]$ is even bounded from below by a positive term. To simplify, let us consider that the observable $\Phi(\Gamma)$ does not include 00, whose condition is met for several problem settings. For instance, we may consider binary classification using trajectories of stochasitc processes. In this case, the observable $\Phi(\Gamma)$ does not include 00. $\Phi(\Gamma)=0$ should be handled separately because $\Phi(\Gamma)=\Phi(\Gamma^{\dagger})=0$ from Eq. (9), showing that the observable is invariant under the time reversal. Later, we will consider the case where $\Phi(\Gamma)=0$ is included. Specifically, it can be proved that the following relation holds:
$C[P(\Phi),P(-\Phi)]\geq\ln 2+H[P(|\Phi|)],$ (11)
whose derivation is shown in Appendix A. In Eq. (11), $P(|\Phi|)$ is the probability distribution of $|\Phi|$, which is the absolute random variable of $\Phi$. Let $\Lambda[P(\Phi)]$ be
$\Lambda[P(\Phi)]\equiv H[P(\Phi)]-H[P(|\Phi|)],$ (12)
which is the entropy difference between the original distribution $P(\Phi)$ and its absolute valued distribution $P(|\Phi|)$. Here, we call $\Lambda[P(\Phi)]$ as symmetry entropy. The symmetry entropy $\Lambda[P(\Phi)]$ quantifies the extent of symmetry of $P(\Phi)$. Figure 1 depicts examples of values of the symmetry entropy, where the horizontal axes are values of $\Phi$ (left column) and $|\Phi|$ (right column) and the vertical axes are $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column). From the condition of time reversal [Eq. (9)], if $\Phi=a$ exists ($a>0$), then $\Phi=-a$ also exists, which is regarded as a pair of the observable. Suppose that the distribution $P(\Phi)$ is symmetric for all of the pairs, as depicted in Fig. 1(a). For this case, $P(|\Phi|)$ is very different from the original distribution $P(\Phi)$, which results in $\Lambda[P(\Phi)]$ being $\ln 2$. In contrast, for an asymmetric distribution illustrated in Fig. 1(b), $P(\Phi)$ and $P(|\Phi|)$ are the same, and thus $\Lambda[P(\Phi)]$ reduces to 00. To be more specific, the symmetry entropy quantifies how biased the probability distributions of pairs of observables are. When there is a strong bias in the probabilities of the pairs, the value of $\Lambda[P(\Phi)]$ becomes small. For example, $\Lambda[P(\Phi)]$ becomes 00 not only in cases where $P(\Phi)$ and $P(|\Phi|)$ are the same, as shown in Fig. 1(b), but also in cases like Fig. 1(c). It can be shown that
$0\leq\Lambda[P(\Phi)]\leq\ln 2,$ (13)
whose proof is provided in Appendix B. In Eq. (13), $\Lambda[P(\Phi)]$ being 00 and $\ln 2$ corresponds to asymmetric and symmetric distributions, respectively. Substituting Eq. (11) into Eq. (10), we obtain the trade-off between the entropy production and the asymmetry of $P(\Phi)$ quantified by entropy:
$\Sigma\geq\ln 2-\Lambda[P(\Phi)]\geq 0.$ (14)
Equation (14) is the main result of this study and referred to as thermodynamic entropic uncertainty relation. The right-hand side of Eq. (14) quantifies the asymmetry of probability distribution $P(\Phi)$. Therefore, Eq. (14) shows that, for arbitrary observable $\Phi$ satisfying the time reversal condition [Eq. (9)], the system requires the entropy production no less than $\ln 2-\Lambda[P(\Phi)]$. The trade-off between entropy production and observable asymmetry parallels traditional thermodynamic uncertainty relations, which illustrate a trade-off between the variance of observables and entropy production. The thermodynamic uncertainty relations consider the relative variance $\mathrm{Var}[\Phi]/\mathbb{E}[\Phi]^{2}$. In a sense, using the relative variance can also be seen as quantifying the asymmetry of $P(\Phi)$; when the variance is smaller and the expectation is greater, the probability distribution $P(\Phi)$ is more asymmetric with respect to $\Phi=0$. Since $\ln 2-\Lambda[P(\Phi)]\geq 0$ in Eq. (14) due to Eq. (13), Eq. (14) can be regarded as a refinement of the second law using the entropy of the observable. There are some advantages in employing the entropy instead of the variance. When dealing with the variance, the observable must yield a real number. However, in cases where the observable consists of classifications such as “success” and “failure”, the variance is not suitable. Even in such cases, the Shannon entropy is well defined, indicating that Eq. (14) can be applied.
Thus far, we assumed that $\Phi(\Gamma)$ takes values in a countable set, that is, $P(\Phi)$ is a discrete distribution. Consider the case where $\Phi(\Gamma)$ produces continuous values, where the summation should be replaced by the integration. When considering a continuous distribution, the notable difference is that the differential entropy may take negative values. However, the existence of negative values is not problematic, as such negative values are offset in $H[P(\Phi)]-H[P(|\Phi|)]$. Therefore, the main result of Eq. (14) is well defined for the continuous case as well.
The derivation of Eq. (14) assumed that the observable $\Phi(\Gamma)$ does not include $\Phi(\Gamma)=0$. It is straightforward to extend the result to the case where the observable includes $\Phi(\Gamma)=0$. Specifically, when $\Phi(\Gamma)=0$ is included, the range of $\Lambda[P(\Phi)]$ is modified as follows:
$0\leq\Lambda[P(\Phi)]\leq[1-P(\Phi=0)]\ln 2$ (15)
The bound becomes
$\Sigma\geq[1-P(\Phi=0)]\ln 2-\Lambda[P(\Phi)]\geq 0,$ (16)
which includes Eq. (14) as the specific case $P(\Phi=0)=0$. As long as the entropy production is given by Eq. (8), Eq. (16) holds for arbitrary observable $\Phi(\Gamma)$ satisfying Eq. (9). In Eq. (16), $\Phi=0$ plays a special role. When considering the process of doing nothing, the observable $\Phi$ is always 00 implying $P(\Phi=0)=1$. For such empty dynamics, the entropy production is 00, showing that both sides of Eq. (16) equal 00 and thus the inequality becomes equality. This consideration shows that $\Phi=0$ should be handled separately. However, note that this exceptional handling of $\Phi=0$ is not limited to Eq. (16). In the conventional thermodynamic uncertainty relation [Eq. (3)], when the expectation of the current vanishes, the inequality is ill-defined. When $P(\Phi)$ is a probability density that is smooth around 00, the measure becomes $P(\Phi=0)=0$. Only when $P(\Phi)$ includes the delta-peaked contribution at $\Phi=0$, we use Eq. (16)
Let us comment on the relation between Eq. (16) and the Landauer principle [30]. When considering a process that resets to one state from a state that exists with equal probability in two states, the following relation is obtained.
$\Delta S_{m}\geq\ln 2,$ (17)
where $\Delta S_{m}$ is the entropy increase in the surrounding medium. Note that the Landauer principle can be considered as a specific case of the second law of thermodynamics. Given the assumption that the system is in steady state, the change in system entropy $\Delta S$ is zero; consequently, $\Sigma=\Delta S_{m}$, where $\Delta S_{m}$ represents the increase in entropy in the environment. Then, Eq. (14) can be expressed by $\Delta S_{m}\geq\ln 2-\Lambda[P(\Phi)]$. When the probability distribution of $P(\Phi)$ ends up with a totally asymmetric distribution, i.e., $\Lambda[P(\Phi)]=0$, Eq. (14) is formally identical to the Landauer principle. However, note that the right side of Eq. (17) arises from a reduction in the Shannon entropy within the system’s state, implying that the Landauer principle is relevant when the system state changes over the time evolution. This contrast with Eq. (14), where the system is assumed to be steady state, and $\ln 2$ on its right-hand side arises from the asymmetry of the observable, not the system state.
Example.—To illustrate Eq. (14), consider a one-dimensional biased random walk as shown in Fig. 2(a). We consider a discrete time model, where the transition probabilities for moving right and left are denoted as $w_{+}$ and $w_{-}$, respectively, with the constraint $w_{+}+w_{-}=1$. The biased random walk plays an important role in biological modeling. For example, a biased random walk can model molecular motors, such as F1-ATPase [31] (Fig. 2(b)). F1-ATPase is an enzyme that can rotate in discrete steps of $32\pi/$ 3 radians in the clockwise or counterclockwise direction [32]. When rotating counterclockwise, F1-ATPase hydrolyzes ATP, while when rotating clockwise, F1-ATPase synthesizes ATP. The rotation in both directions is stochastic, which means that it occurs randomly. The counter-clockwise rotation can be considered as moving in the positive direction in the biased random walk, where the motor has a higher probability of rotating in the counter-clockwise direction.
In the biased random walk, we consider the observable $\Phi(\Gamma)$, which takes either “$+$” or “--”. After the $L$ steps, the observable $\Phi$ is determined by comparing the final position to the initial position. If the final position is greater than the initial position, $\Phi$ is assigned the value “$+$”. Conversely, if the final position is less than or equal to the initial position, $\Phi$ is assigned the value “--”. When the number of steps is odd, there is no possibility of $\Phi(\Gamma)=0$. When considering the aforementioned molecular motor, this observable corresponds to checking whether the motor rotates in the counterclockwise direction or not. For details, see Appendix C.
Figure 3: Results of the biased random walk. The sum of the entropy production and the symmetry entropy, $\Sigma+\Lambda[P(\Phi)]$, is plotted as a function of $w_{+}$ with the solid line. The dotted line denotes $\ln 2$, which is the lower bound of $\Sigma+\Lambda[P(\Phi)]$ in Eq. (14). Each of these terms, $\Sigma$ and $\Lambda[P(\Phi)]$, is plotted by the dashed line and the dot-dashed line, respectively. Here, we employ $L=5$.
We calculate quantities, such as $\Sigma$ and $\Lambda[P(\Phi)]$, as a function of $w_{+}$ in Fig. 3. Note that, for the binary case where $\Phi(\Gamma)$ takes two values, $\Lambda[P(\Phi)]=H[P(\Phi)]$. We consider the steady-state condition and calculate the probability distribution $P(\Phi)$ and $\Sigma$. In Fig. 3, $\Sigma+\Lambda[P(\Phi)]$, the sum of the entropy production $\Sigma$ and the symmetry entropy $\Lambda[P(\Phi)]$, is represented by the solid line, where the dotted line indicates $\ln 2$. As can be seen in the figure, $\Sigma+\Lambda[P(\Phi)]$ is always above $\ln 2$, confirming the validity of Eq. (14). To examine how each of these components, $\Sigma$ and $\Lambda[P(\Phi)]$, behave as a function of $w_{+}$, $\Sigma$ and $\Lambda[P(\Phi)]$ are represented by the dashed line and the dot-dashed line, respectively. Regarding entropy production $\Sigma$, it falls below $\ln 2$ in the region where $w_{+}$ ranges from $0.5$ to $0.65$. On the other hand, the symmetry entropy of the observable $\Lambda[P(\Phi)]$ is decreasing as a function of $w_{+}$, and except for $w_{+}=0.5$, it is below $\ln 2$. From these results, we can conclude that the inequality $\Sigma+\Lambda[P(\Phi)]\geq\ln 2$ holds only when considering the sum, $\Sigma+\Lambda[P(\Phi)]$.
Conclusion.—In this study, we established a thermodynamic entropic uncertainty relation that links entropy production and the Shannon entropy of the observables. Our findings extend conventional thermodynamic uncertainty relations by incorporating measures based on entropy, highlighting the role of Shannon entropy of observables in stochastic thermodynamics. The derived inequality formalizes a fundamental trade-off between entropy production and the asymmetry of the observable distribution. We validated this relation using a biased random walk model. This framework provides a deeper understanding of nonequilibrium thermodynamics and expands the application of entropy-based uncertainty relations in stochastic systems. One direction of expansion is towards quantum systems. In recent years, the thermodynamic uncertainty relations in quantum systems [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 14, 43] have garnered significant attention. In particular, the thermodynamic uncertainty relations within the framework of continuous measurement are closely related to those in classical stochastic processes. This direction presents future challenges.
Appendix A Derivation of the main result [Eq. (16)]
A.1 Discrete case
In this section, we derive Eq. (16). Let us introduce the absolute random variable of $\Phi$, which is expressed by $|\Phi|$. The probability distribution of $|\Phi|$ is given by
$P(|\Phi|=\phi)=\begin{cases}P(\Phi=\phi)+P(\Phi=-\phi)&\phi\neq 0\\ P(\Phi=0)&\phi=0\end{cases}.$ (18)
Therefore, the Shannon entropy of $|\Phi|$ is given by
$H[P(|\Phi|)]$
$=-\sum_{\phi\geq 0}P(|\Phi|=\phi)\ln P(|\Phi|=\phi)$
$=-P(0)\ln P(0)-\sum_{\phi>0}[P(\phi)+P(-\phi)]\ln[P(\phi)+P(-\phi)]$
$=-P(0)\ln P(0)-\sum_{\phi>0}P(\phi)\ln[P(\phi)+P(-\phi)]$
$-\sum_{\phi<0}P(\phi)\ln[P(\phi)+P(-\phi)]$
$=-\sum_{\phi}P(\phi)\ln[P(\phi)+P(-\phi)]+P(0)\ln 2,$ (19)
where we abbreviated $P(\phi)=P(\Phi=\phi)$. Then we compute $H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$ as follows:
$H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$
$=\sum_{\phi}P(\phi)\ln\frac{P(-\phi)/P(\phi)}{1+P(-\phi)/P(\phi)}+P(0)\ln 2.$ (20)
Let us consider the function $f(x)=\ln\frac{x}{1+x}$. Since $f(x)$ is concave for $x>0$, by using the Jensen inequality, the following relation holds:
$\sum_{\phi}P(\phi)\ln\frac{P(-\phi)/P(\phi)}{1+P(-\phi)/P(\phi)}$
$\leq\ln\frac{\sum_{\phi}P(\phi)\left(P(-\phi)/P(\phi)\right)}{1+\sum_{\phi}P(\phi)\left(P(-\phi)/P(\phi)\right)}$
$=-\ln 2.$ (21)
By substituting Eq. (21) into Eq. (20), we obtain
$H[P(|\Phi|)]+[1-P(\Phi=0)]\ln 2$
$\leq C[P(\Phi),P(-\Phi)].$ (22)
Equation (22) proves Eq. (11) in the main text.
A.2 Continuous case
Next, we consider the continuous case. Basically, the derivation is the same as in the discrete case except that the summation should be replaced by the integration.
Assume $P(\Phi)$ represents a probability density that is differentiable for all $\Phi$. For the continuous case, the probability density $P(|\Phi|)$ is defined by
$P(|\Phi|=\phi)=P(\Phi=\phi)+P(\Phi=-\phi).$ (23)
We do not consider $P(|\Phi|=0)$, because such measure is 00 for the smooth probability density. When $P(\Phi)$ includes the contribution of the delta function in $\Phi=0$, this is not the case. The continuous entropy is given by
$H[P(|\Phi|)]$
[…]
Again, we calculate $H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$ as follows:
$H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$
[…]
Using Eq. (25), we prove
$H[P(|\Phi|)]+\ln 2$ $\leq C[P(\Phi),P(-\Phi)].$ (26)
Appendix B Proof of Eq. (13)
In this section, we prove Eq. (13). Here, we show the relation for $P(\Phi=0)\geq 0$. The relation which we want to show is given by
$0\leq\Lambda[P(\Phi)]\leq[1-P(\Phi=0)]\ln 2$ (27)
The first inequality part corresponds to $H[P(|\Phi|)]\leq H[P(\Phi)]$. For random variables $X$ and $Y$, the following relation holds:
$H[P(X),P(Y)]$ $=H[P(X)|P(Y)]+H[P(Y)]$
$=H[P(Y)|P(X)]+H[P(X)].$ (28)
Substituting $X=\Phi$ and $Y=|\Phi|$, we obtain
$H[P(|\Phi|)|P(\Phi)]+H[P(\Phi)]$ $=H[P(\Phi)]$
$=H[P(\Phi)|P(|\Phi|)]+H[P(|\Phi|)]$
$\geq H[P(|\Phi|)],$ (29)
where we used $H[P(|\Phi|)|P(\Phi)]=0$. Equation (29) proves the first inequality part of Eq. (27). Note that, for the continuous case, this proof does not work, as the continuous conditional entropy $H[P(\Phi)|P(|\Phi|)]$ may take negative values. However, we can show that the first inequality part also holds for the continuous case by a direct calculation.
Next, we prove the second part of the inequality of Eq. (27), which can be done following the same approach as in Appendix A. Using Eq. (19), we have
$H[P(|\Phi|)]-H[P(\Phi)]$
$=-\sum_{\phi}P(\phi)\ln\left(1+\frac{P(-\phi)}{P(\phi)}\right)+P(0)\ln 2$
$\geq-\ln 2+P(0)\ln 2,$ (30)
where we again used the Jensen inequality. Equation (30) proves the second inequality part of Eq. (27). For the continuous case, by following the same argument by replacing the summation with the integration, the following relation holds:
$H[P(|\Phi|)]-H[P(\Phi)]$ $\geq-\ln 2.$ (31)
Appendix C Details of example
We show details of the biased random walk example. The probability distribution at position $x$ after $L$ steps is given by
$p(x)=\begin{cases}{\displaystyle\binom{L}{\frac{L+x}{2}}p^{\frac{L+x}{2}}(1-p)^{\frac{L-x}{2}}}&L+x\in\mathrm{even}\\ 0&\mathrm{otherwise}\end{cases}.$ (32)
Using $p(x)$ in Eq. (32), for odd $L$, the probability of $\Phi$ is calculated by $P(\Phi=\mathrm{decrease})=\sum_{x=-L}^{-1}p(x)$ and $P(\Phi=\mathrm{increase})=\sum_{x=1}^{L}p(x)$, . which can be used to calculate the Shannon entropy of $P(\Phi)$. Moreover, under the steady-state condition, the entropy production $\Sigma$ up to $L$ steps is given by
$\Sigma=L(w_{+}-w_{-})\ln\frac{w_{+}}{w_{-}}.$ (33)